Exercise 31.33 Part A Youre an optometrist helping a nearsighted patient who claims he cant see clearly beyond 70.0 cm . Prescribe a lens that will put the images of distant objects 70.0 cm , giving your patient clear vision at all distances beyond the normal near point. Express your answer using two significant figures. diopters Submit My Answers Give Up Provide Feedback Continue
Answer #1


object distance p =infinity


image distance q = -70 cm

focal length f = ?

1/p + 1/q = 1/f


1/infinity - 1/70 = 1/f

focal langth f = -70 cm = -0.7 m


power of lens P = 1/f = 1/0.7 = -1.43 diopters